A fair coin is tossed 5 successive times . what is the probability that at least one head is obtained ?
a) 8/10 b) 9/10 c) 1 d) 31/32 e)none

Answer by stanbon(75887)

A fair coin is tossed 5 successive times . what is the probability that at least one head is obtained ?
----
P(at least one) = 1 - P(none)
-----
P(at least one) = 1 - (1/2)^5
= 1 - 1/32 = 31/32
=========================
Cheers,
Stan H.

a) 8/10 b) 9/10 c) 1 d) 31/32 e)none